OpenStudy (anonymous):

write 6% as a decimal

OpenStudy (unklerhaukus):

\[6\%\] \[6\%=\frac{6}{100}=\frac{0.6}{10}=\frac{0.06}{1}=0.06\]

OpenStudy (anonymous):

to convert to decimal % sign means divide by 100 as 6/100=0.06

OpenStudy (anonymous):

Thank you no matter what I do I am no good at this stuff, being out of school 20yrs and jumping bck into it has made me a nervous wreck